mathslover (mathslover):

Hello everyone , I need help in this one : \[\large{\frac{2^x}{1+2^x}=\frac{1}{4}}\] , then the value of \[\large{\frac{8^x}{1+8^x}}\]

OpenStudy (lgbasallote):

is that second part related to the first part?

mathslover (mathslover):

I am getting upto this step : \(\large{2^{x+2}=1+2^x}\)

mathslover (mathslover):

yes may be lemme try that also @lgbasallote

mathslover (mathslover):

may I say something? : I am not getting the one that I should have..

mathslover (mathslover):

I got this... may be I made this more complex..

OpenStudy (lgbasallote):

what if in the first part you let a = 2^x \[\implies \frac{a}{1+a} = \frac 14\] cross multiply \[\implies 4a = 1 + a\] subtract a from both sides \[\implies 3a = 1\] divide both sides by 3 \[\implies a = \frac 13\] sub back a = 2^x \[\implies 2^x = \frac 13\] use log \[\implies \log_2 (\frac 13) = x\] does that help?

mathslover (mathslover):

is there any involvement of log.. in the first expression? (eqn)

mathslover (mathslover):

hmn right I will use this wait

OpenStudy (kinggeorge):

If you didn't see that method, you can also go this way.\[2^{x+2}-2^x=1\]\[2^x(4-1)=1\]\[2^x=\frac{1}{3}\]

mathslover (mathslover):

right I am doing that .. kg sir

mathslover (mathslover):

Oh k so i got \(\frac{1}{28}\)

mathslover (mathslover):

it's in option and is right.. also the process is excellent thanks to both @lgbasallote and @KingGeorge

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (lgbasallote):

oh..cool...it was a questionception

mathslover (mathslover):

:)

mathslover (mathslover):

btw can any1 tell me that these type of questions come into which topic?

mathslover (mathslover):

Basic algebra or exponents ? something like that